OpenStudy (meilendurcer):

sqrt 10 multiplied by sqrt 8 = sqrt 80?

OpenStudy (meilendurcer):

@hello1213 please help.

OpenStudy (mathmale):

Let's write this out symbolically:\[\sqrt{10}*\sqrt{8}=\sqrt{8*10}?\]What do you think?

OpenStudy (anonymous):

looks good to me

OpenStudy (anonymous):

on the other hand, since \(80=16\times 5\) this is the same as \[\sqrt{80}=\sqrt{16\times 5}=\sqrt{16}\sqrt{5}=4\sqrt5\]

OpenStudy (meilendurcer):

ahh good that was my second choice. thanks!

OpenStudy (anonymous):

yw